how to I find the value of the legs if I have the value of the hypotenuse? the hypotenuse = 10m but we don't have the value of the legs!! thank you!!
Found 2 solutions by rothauserc, ikleyn:
Answer by rothauserc(4718)

a^2 + b^2 = 10^2
:
a^2 + b^2 = 100
:
64 + 36 = 100
:
8^2 + 6^2 = 10^2
:
************************
one possible solution is
8 and 6 for the legs
************************
:

Answer by ikleyn(52781)

.
If you are given the hypotenuse length only, you CAN NOT determine the lengths of the legs.

Too few data. You need an additional info.